# Dormant BH candidates from Gaia DR3 summary diagnostics

### v1

@@ 5.1 Contamination of the Final Candidate Sample
Several sources of contamination will affect our RGB+BH/NS candidate sample. Here we summarise the most relevant effects (see Sect. 4.4 for details on the applied quality cuts).
Our inference is based on Gaia DR3 summary statistics, particularly ruwe and $\sigma_{\mathrm{RV}}$. While both are sensitive to binarity, they can also be affected by other circumstances. For example, marginally resolved wide binaries can distort the astrometric solution and increase astrometric excess noise (El-Badry et al., 2024a), while stellar activity, such as pulsations or star spots, can inflate RV scatter. To mitigate these, we excluded stars with variability inconsistent with ellipsoidal modulation, as well as obvious eclipsing binaries (Sect. 4.4 C).
Underestimated parallaxes may also cause sample contaminants. To address this, we removed stars with strongly discrepant Gaia parallaxes relative to spectrophotometric estimates (Zhang et al., 2023, Sect. 4.4 B). However, moderate parallax underestimates in the remaining sample may still lead to overestimated companion masses.
Companion masses can also be overestimated for binaries with nearly edge-on inclinations, for which the inferred companion mass is systematically inflated when marginalised over an assumed isotropic inclination distribution. This introduces contamination from RGB binaries with main-sequence or with white dwarf companions (see Sect. 5.2).
@@
Contamination from hierarchical triple systems is another major concern (Andrew et al., 2022). In such systems, a short-period inner binary can enhance RV variability, while a distant tertiary induces astrometric motion, jointly mimicking the signal of a single massive companion. These configurations can lead to spurious inferences of an intermediate-period binary with a compact object. Similar contamination has been identified in earlier searches for dark companions based on high ruwe and RV variability (Andrew et al., 2022). To minimise this, we restricted our analysis to evolved stars, for which short-period inner binaries should not exist because of the large stellar radii. However, triples with moderately separated components (e.g. 200-day and 2000-day orbits) could still remain in the sample.
@@
Despite the cleaning steps applied, it is expected that a large fraction of contamination remains, requiring spectroscopic follow-up to identify and eliminate false positives more robustly (see Sect. 5.3).
5.2 Comparison with Gaia orbital solutions
Figure 10: Comparison between binary parameters inferred from Gaia summary diagnostics and reference values from the DR3 NSS catalogue. The left panels show histograms for the 10,298 stars (out of 21 thousand variable RGB stars) with Gaia orbital solutions, while the three right panels show the subsets with AstroSpectroSB1, SB1, and Orbital solutions. Comparisons are shown for orbital periods (top) and companion masses (bottom). Gaia BH 2 and the candidate BH from LAMOST are highlighted (black). The systematic offsets in the inferred parameters likely reflect a combination of inclination effects, the preferentially high inclinations and shorter periods of systems in the NSS sample, and the fact that spectroscopic orbits yield only minimum companion masses.
We address two relevant questions regarding the Gaia NSS orbital catalogue; (a) For RGB+BH candidates that also have NSS orbital solutions, how well do the published orbits agree with our inferred parameters? (b) Given the inferred parameters, how many of our candidates would be expected to receive orbital solutions in DR3, and what fraction are likely to receive orbital solutions in the upcoming DR4 release?
For part a) we compared inferred periods and companion masses to published values from the NSS catalogue. Of the 21,028 RGB stars that passed the initial cuts on excess scatter in RV and position, 10,712 have NSS binary solutions. After applying the cleaning steps from Sect. 4.4 (removing photometric variables, unreliable parallaxes, and $\sigma_{\mathrm{RV}}$ outliers), the number reduces to 19,887 stars of which 10,298 have NSS solutions. [^1] Among these 19,887 stars, 1006 have inferred companion masses exceeding $3\,\mathrm{M}_{\odot}$, of which 450 also have published NSS solutions. After removing stars with NSS solutions, 556 high-mass companion candidates remain in our final catalogue (Sect. 4.4).
Figure 10 (top) shows the comparison of catalogued vs. inferred orbital periods for these 10,298 objects (similar to Andrew et al., 2022, their Fig. 9). Despite the limited number of observables and many degrees of freedom in the inference problem, we find a good overall correspondence. Some scatter is expected, given the large uncertainties and typically broad posterior distributions (see Fig. 3). At long periods ($\sim$1000 d), our method tends to overestimate the orbital period. This probably reflects selection effects in the NSS catalogue: binaries with $P\gtrsim 1000$ d are more often given astrometric acceleration or spectroscopic trend solutions rather than full orbital solutions, so the NSS comparison sample is biased toward shorter periods.
We also compare the inferred companion masses with the NSS values (Fig. 10, bottom). Again, a linear relation is evident, though our method systematically yields higher masses by an average of 39%. Several effects contribute here, mainly related to orbital inclination. First, NSS spectroscopic solutions lack inclination constraints, so their companion masses represent lower limits. Second, the sample of stars with spectroscopic solutions is biased toward edge-on systems, where RV amplitudes are the largest. In contrast, in our forward modelling we assumed an isotropic distribution of inclinations (uniform in $\cos i$). This marginalisation overestimates masses for edge-on configurations; if these are more prevalent in the comparison sample, it causes a systematic offset. An additional contribution may come from our initial selection cut of rv_amplitude_robust >20>20\,km s-1, which biases the sample toward high-rv_amplitude_robust and $\sigma_{\mathrm{RV}}$ outliers. Since massive companions are intrinsically rare, this selection likely favours systems with overestimated amplitudes due to noise or sampling effects. Finally, for some NSS systems, the uncertainties in inclination also add scatter, meaning that the uncertainties on the $x$-axis can also be large.
@@ 5.2 Comparison with Gaia orbital solutions
For part (b), we followed the approach used in Sect. 4.5 and forward-modelled the Gaia astrometric processing pipeline with gaiamock. For each of the 556 RGB+BH candidates in the cleaned catalogue, we generated 20 realisations of the binary parameters as random draws from the weighted posterior samples. To account for underestimated parallax errors in high-ruwe stars, parallaxes were resampled with inflated uncertainties following El-Badry (2025). Each realisation was then passed through the simulated Gaia astrometric processing chain to estimate the probability of a given solution type (single-star 5-parameter, acceleration 7/9-parameter, or 12-parameter orbit).
@@
Summing over all candidates, we find that $\sim$81% are expected to receive single-star solutions in DR3, 4.6% acceleration solutions, and only 14.7% full 12-parameter orbital solutions. The small predicted fraction of orbital solutions is reassuring: if many more of our candidates should already have received 12-parameter solutions, this would cast doubt on the inferred parameters. For individual stars, the probability of an orbit solution is low, with only one candidate exceeding a 50% probability. Under the assumption that the inferred parameters are correct and using the same NSS quality cuts as for DR3, we predict a substantial increase in orbital solutions in DR4. The simulations suggest that about 60% of our candidates could then receive astrometric binary orbits.
@@ 5.3 Spectroscopic follow-up
We have initiated a comprehensive programme of spectroscopic follow-up for the most promising candidates in our sample and are in the process of obtaining observations. Although complete phase coverage may be infeasible because of the typically long orbital periods (on the order of several years), even few-epoch RV data can be useful when combined with the epoch astrometry and RVs expected from Gaia DR4. These data will extend the observing baseline and enable joint orbital fits, improving parameter constraints and the prospects of confirming dormant NS or BH companions. The results of the follow-up will be described in a forthcoming paper.
@@ 6 Summary and conclusions
Most known stellar-mass BHs have been detected in X-ray binaries or through gravitational wave events. These techniques probe only a minute and biased subset of the total population. Detecting larger numbers of non-accreting dormant BHs will enable exploration of bulk properties of the Galactic BH population and help improve our understanding of binary evolution processes and supernova physics.
@@
Spectroscopic follow-up remains essential to validate candidates and eliminate false positives. High-resolution, multi-epoch RV monitoring can confirm the binary origin of variability, identify spurious systems with luminous secondaries or hierarchical triples, and provide independent constraints on orbital parameters and minimum companion mass. Combining follow-up RVs with future Gaia DR4 epoch astrometry will enable joint orbital fits and secure identification of dormant NS/BH companions.
@@
The full catalogue of 21,028 RGB stars with excess RV and astrometric scatter, including inferred orbital periods and companion masses, is available on Zenodo. The data set also contains additional parameters and quality flags that allow users to select the 556 most promising RGB+BH candidates from the cleaned catalogue. An analogous catalogue for 19,664 main-sequence star candidates is also provided and described in Appendix B.
@@ Acknowledgements.
J. Müller-Horn thanks Thibault Merle for helpful discussions and suggestions regarding the use of the SBX binary catalogue. J. Müller-Horn, B. Pennell, M. Green, J. Li, R. Seeburger, and H. W. Rix acknowledge support from the European Research Council for the ERC Advanced Grant [101054731]. This work has made use of data from the European Space Agency (ESA) mission Gaia (https://www.cosmos.esa.int/Gaia), processed by the Gaia Data Processing and Analysis Consortium (DPAC, https://www.cosmos.esa.int/web/Gaia/dpac/consortium). Funding for the DPAC has been provided by national institutions, in particular the institutions participating in the Gaia Multilateral Agreement.
@@ Appendix A Stellar parameter estimation with XGBoost
Figure 11: Performance of the XGBoost model in predicting stellar masses and radii for RGB stars. The top and bottom panels show the true vs. predicted values for mass and radius, respectively, based on the validation set. Reference values are asteroseismic measurements from the APOKASC-3 catalogue. The model achieves typical fractional uncertainties of $\sim$15% for mass and $\sim$7% for radius, with small systematic bias.
@@
The model achieved mean cross-validation RMSEs of $0.21\pm 0.01\,\text{M}_{\odot}$ for stellar mass and $1.23\pm 0.15\,\text{M}_{\odot}$ for stellar radius. The predicted vs. true values for the test set are shown in Fig. 11. For stellar masses, we determined a mean and standard deviation of the logarithmic ratio $\log(M_{\mathrm{pred}}/M_{\mathrm{true}})=0.011\pm 0.150$, corresponding to a typical fractional uncertainty of $\sim$15% and minimal bias. For stellar radii, we obtained $\log(R_{\mathrm{pred}}/R_{\mathrm{true}})=0.005\pm 0.069$, or $\sim$7% uncertainty. As can be expected, the effective temperature and the absolute magnitude (in W1) were the most important training features.
@@ Appendix B A catalogue of dormant BH candidates on the main sequence
The main challenge for MS stars is the increased contamination from luminous binaries and triples (Andrew et al. 2022). To mitigate this, we used the data-driven MS+MS binary catalogue of Li et al. (2025), based on composite BP/RP spectral signatures, to exclude systems with clearly luminous companions.
@@
The first six cuts select MS stars with $M_{G}>3$ mag for which the BP/RP filtering of luminous binaries is applicable (Li et al. 2025), while the last four cuts match those applied to the RGB sample (Sect. 4), selecting systems with excess scatter in RV and astrometry.
@@
We then applied the same forward-modelling inference as for the RGB sample, fitting companion masses and orbital periods for all 19,664 MS stars. Figure 13 shows the inferred distributions. A total of 711 systems have companion masses above 3 M. After applying the quality cuts described in Sect. 4.4, that is, removing 225 stars with spurious RV scatter, one star with an unreliable parallax, 163 photometrically variable stars, and 43 with NSS orbital solutions, the final cleaned catalogue contains 279 MS+BH candidates. Their CMD distribution is shown in Fig. 14. The apparent clustering of candidates at $M_{G}\lesssim 4.5$ mag probably reflects residual contamination from luminous binaries. For bright, upper-MS stars, age effects complicate isochrone fitting, and BP/RP spectral decomposition becomes less robust (Li et al. 2025). In contrast, the small number of promising candidates at fainter magnitudes (only seven candidates with $M_{G}\geq 4.5$ mag) makes spectroscopic follow-up both valuable and feasible in this regime.
@@
Figure 14: CMD of the 279 final MS+BH candidates (white stars), overlaid on the initial MS candidate sample (2D histogram). The confirmed BH binary Gaia BH 1 is highlighted, together with known Gaia NS binaries (triangles). The apparent clustering of candidates at $M_{G}<4.5$ mag (shaded grey) likely reflects the less efficient removal of MS+MS binaries on the upper main sequence by the BP/RP-based filtering.
Of the known compact-object binaries with MS primaries (including Gaia BH 1 and several MS+NS systems), only two are included in our parent sample of 19,664 stars. The others are excluded because they lack epoch RV measurements (rv_method_used $\neq 1$) and therefore do not meet the initial selection criteria. For the two MS+NS systems that are included (Gaia DR3 5136025521527939072 and 1801110822095134848, with periods of 537 and 894 days, respectively; El-Badry et al. 2024b), our fits recover long periods ($P_{\mathrm{fit}}\approx 600$ and 800 d) and high companion masses ($M_{\mathrm{CO,fit}}\approx 2.4$ and 2.9 M). As expected for NS companions, these fall below the $3\,M_{\odot}$ BH threshold, but would be retained if the threshold were lowered slightly.
To further validate the method, we compared our inferred periods and masses to Gaia DR3 NSS solutions for the 10,298 MS stars in the parent sample with AstroSpectroSB1, SB1, or Orbital entries. Figure 15 shows that our fits reproduce the NSS parameters with good correspondence, despite being based only on two observables (ruwe and $\sigma_{\mathrm{RV}}$). However, for very short periods ($\lesssim 50$ d), we often overpredict the orbital period. This bias likely arises from contamination by hierarchical systems, where short-period inner binaries inflate the RV scatter while long-period tertiaries increase ruwe values, leading our inference framework to favour intermediate periods.
Figure 15: Comparison between binary parameters inferred from Gaia summary diagnostics and reference values from the DR3 NSS catalogue for the MS sample. Shown are histograms for the 6195 stars (out of 19,664 fitted MS stars) with AstroSpectroSB1, SB1, or Orbital NSS solutions. Panels compare orbital periods (top) and companion masses (bottom). The two known Gaia NS binaries in this sample are highlighted in black. Despite relying only on two observables (ruwe and $\sigma_{\mathrm{RV}}$), our fits reproduce the catalogued parameters with good overall correspondence.

### v2

@@ 5.1 Contamination of the final candidate sample
Several sources of contamination will affect our RGB+BH candidate sample. Here we summarise the most relevant effects (see Sect. 4.4 for details on the applied quality cuts).
Our inference is based on Gaia DR3 summary statistics, particularly ruwe and $\sigma_{\mathrm{RV}}$. While both are sensitive to binarity, they can also be affected by other circumstances. For example, marginally resolved wide binaries can distort the astrometric solution and increase astrometric excess noise (Gaia Collaboration et al., 2023a; Halbwachs et al., 2023; Holl et al., 2023; El-Badry et al., 2024a; Gosset et al., 2025). This effect motivated the stringent parallax signal-to-noise cuts applied in DR3 for astrometric binary solutions. Furthermore, stellar activity, such as pulsations or star spots, can inflate RV scatter. To mitigate contamination from these non-orbital effects, we excluded stars with variability inconsistent with ellipsoidal modulation, as well as obvious eclipsing binaries (Sect. 4.4 C).
Erroneous parallaxes may also cause sample contaminants, as the inferred companion mass depends on the assumed physical scale of the orbit. In our inference, the parallax was treated as a free parameter, so that the uncertainties in $\varpi$ were consistently propagated in the inferred orbital parameters and companion masses. To address extreme cases, we further removed stars with strongly discrepant Gaia parallaxes relative to spectrophotometric estimates (Zhang et al., 2023, Sect. 4.4 B). However, moderate parallax underestimates in the remaining sample may still lead to overestimated companion masses.
Companion masses can also be overestimated for binaries with nearly edge-on inclinations, for which the inferred companion mass is systematically inflated when marginalised over an assumed isotropic inclination distribution. This introduces contamination from RGB binaries with companions on the main sequence or with white dwarf companions (see Sect. 4.5).
@@
Contamination from hierarchical triple systems is another major concern (Andrew et al., 2022). In such systems, a short-period inner binary can enhance RV variability, while a distant tertiary induces astrometric motion, jointly mimicking the signal of a single massive companion. These configurations can lead to spurious inferences of an intermediate-period binary with a compact object. Similar contamination has been identified in earlier searches for dark companions based on high ruwe and RV variability (Andrew et al., 2022). To minimise this, we restricted our analysis to evolved stars; while stars low on the RGB may still host short-period inner binaries, the large current or past radii of more evolved giants preclude such systems. However, triples with moderately separated components (e.g. 200-day and 2000-day orbits) could still remain in the sample.
@@
Despite the cleaning steps applied, it is expected that a high fraction of contamination remains, requiring spectroscopic follow-up to identify and eliminate false positives more robustly (see Sect. 5.3).
5.2 Expected Gaia orbital solutions for RGB+BH candidates
@@
As a consistency check on the inferred binary parameters, we estimated how many of the RGB+BH candidates would be expected to receive orbital solutions in Gaia DR3 and which fraction are likely to do so in the upcoming DR4 release. We followed the approach used in Sect. 4.6 and forward modelled the Gaia astrometric processing pipeline with gaiamock. For each of the 389 RGB+BH candidates in the cleaned catalogue, we generated 20 realisations of the binary parameters as random draws from the weighted posterior samples. Each realisation was then passed through the simulated Gaia astrometric processing chain to estimate the probability of a given solution type (single-star 5-parameter, acceleration 7- or 9-parameter, or 12-parameter orbit).
@@
Summing over all candidates, we find that $\sim$74.4% are expected to receive single-star solutions in DR3, 5.9% acceleration solutions, and only 19.8% full 12-parameter orbital solutions. The moderate predicted fraction of orbital solutions is reassuring: if many more of our candidates should already have received 12-parameter solutions, this would cast doubt on the inferred parameters. For individual stars, the probability of an orbit solution is low, with only one candidate exceeding a 50% probability. Under the assumption that the inferred parameters are correct and using the same NSS quality cuts as for DR3, we predict a substantial increase in orbital solutions in DR4. The simulations suggest that about 63% of our candidates could then receive astrometric binary orbits.
@@ 5.3 Spectroscopic follow-up
We have begun an observing campaign for spectroscopic follow-up of the most promising candidates in our sample using the FEROS instrument at the La Silla MPG 2.2 m telescope with approximately 50 nights of allocated observing time. Although complete phase coverage may be infeasible because of the typically long orbital periods (on the order of several years), even few-epoch RV data can be useful when combined with the epoch astrometry and RVs expected from Gaia DR4. These data will extend the observing baseline and enable joint orbital fits, improving parameter constraints and the prospects of confirming dormant NS or BH companions. The results of the follow-up will be described in a forthcoming paper.
@@ 6 Summary and conclusions
Most known stellar-mass BHs have been detected in X-ray binaries or through gravitational wave events. These techniques probe only a minute and biased subset of the total population. Detecting larger numbers of non-accreting dormant BHs will enable the exploration of the bulk properties of the Galactic BH population and help improve our understanding of binary evolution processes and supernova physics.
@@
Spectroscopic follow-up remains essential to validate candidates and eliminate false positives. High-resolution, multi-epoch RV monitoring can confirm the binary origin of variability, identify spurious systems with luminous secondaries or hierarchical triples, and provide independent constraints on orbital parameters and minimum companion mass. Combining follow-up RVs with future Gaia DR4 epoch astrometry will enable joint orbital fits and secure identification of dormant NS and BH companions.
@@
The full catalogue of 21,028 RGB stars with excess RV and astrometric scatter, including inferred orbital periods and companion masses, is available on https://doi.org/10.5281/zenodo.17271785. The dataset also contains additional parameters and quality flags that allow users to select the 389 most promising RGB+BH candidates from the cleaned catalogue. An analogous catalogue for 19,664 main-sequence star candidates is also provided and described in Appendix B.
@@ Acknowledgements.
The authors sincerely thank the anonymous referee for their thorough review and thoughtful and constructive comments. J. Müller-Horn thanks Thibault Merle for helpful discussions and suggestions regarding the use of the SBX binary catalogue. J. Müller-Horn, B. Pennell, M. Green, J. Li, R. Seeburger, and H. W. Rix acknowledge support from the European Research Council for the ERC Advanced Grant [101054731]. This work has made use of data from the European Space Agency (ESA) mission Gaia (https://www.cosmos.esa.int/Gaia), processed by the Gaia Data Processing and Analysis Consortium (DPAC, https://www.cosmos.esa.int/web/Gaia/dpac/consortium). Funding for the DPAC has been provided by national institutions, in particular the institutions participating in the Gaia Multilateral Agreement.
@@ Appendix A Stellar parameter estimation with XGBoost
Figure 11: Performance of the XGBoost model in predicting stellar masses and radii for RGB stars. The top and bottom panels show the true vs predicted values for mass and radius, respectively, based on the validation set. The reference values are asteroseismic measurements from the APOKASC-3 catalogue. The model achieves typical fractional uncertainties of $\sim$15% for mass and $\sim$7% for radius, with small systematic bias.
@@
The model achieved mean cross-validation RMSEs of $0.21\pm 0.01\,\text{M}_{\odot}$ for stellar mass and $1.23\pm 0.15\,\text{M}_{\odot}$ for stellar radius. The predicted vs true values for the test set are shown in Fig. 11. For stellar masses, we determined a mean and standard deviation of the logarithmic ratio $\log(M_{\mathrm{pred}}/M_{\mathrm{true}})=0.011\pm 0.150$, corresponding to a typical fractional uncertainty of $\sim$15% and minimal bias. For stellar radii, we obtained $\log(R_{\mathrm{pred}}/R_{\mathrm{true}})=0.005\pm 0.069$, or $\sim$7% uncertainty. As can be expected, the effective temperature and the absolute magnitude (in W1) were the most important training features. [^3] These absolute magnitudes were computed without applying the Gaia DR3 parallax zero-point correction (Lindegren et al. 2021a), which would slightly reduce the inferred stellar masses of some stars. The effect is generally small compared to the XGBoost fit uncertainties, being less than 0.14 M for more than 99% of RGB candidates.
@@ Appendix B A catalogue of dormant BH candidates on the main sequence
The main challenge for MS stars is the increased contamination from luminous binaries and triples (Andrew et al. 2022). To mitigate this, we used the data-driven MS+MS binary catalogue of Li et al. (2025), based on composite XP spectral signatures, to exclude systems with clearly luminous companions.
@@
The first six cuts select MS stars with $M_{G}>3$ mag for which the XP filtering of luminous binaries is applicable (Li et al. 2025), while the last four cuts match those applied to the RGB sample (Sect. 4), selecting systems with excess scatter in RV and astrometry.
@@
We then applied the same forward-modelling inference as for the RGB sample, fitting companion masses and orbital periods for all 19,664 MS stars. For these stars, we kept the parallax fixed to the DR3 catalogue value, since the sample is limited to nearby stars with high parallax significance ($\varpi/\sigma_{\varpi}>10$) and the Gaia zero-point correction is less relevant at these distances. Figure 13 shows the inferred distributions. A total of 711 systems have companion masses above 3 M. After applying the quality cuts described in Sect. 4.4, that is, removing 225 stars with spurious RV scatter, one star with an unreliable parallax, 163 photometrically variable stars, and 43 with NSS orbital solutions, the final cleaned catalogue contains 279 MS+BH candidates. Their CMD distribution is shown in Fig. 14. The apparent clustering of candidates at $M_{G}\lesssim 4.5$ mag probably reflects residual contamination from luminous binaries. For bright, upper-MS stars, age effects complicate isochrone fitting, and XP spectral decomposition becomes less robust (Li et al. 2025). In contrast, the small number of promising candidates at fainter magnitudes (only seven candidates with $M_{G}\geq 4.5$ mag) makes spectroscopic follow-up both valuable and feasible in this regime.
@@
Figure 14: CMD of the 279 final MS+BH candidates (white stars), overlaid on the initial MS candidate sample (2D histogram). The confirmed BH binary Gaia BH 1 is highlighted, together with known Gaia NS binaries (triangles). The apparent clustering of candidates at $M_{G}<4.5$ mag (shaded grey) likely reflects the less efficient removal of MS+MS binaries on the upper main sequence by the XP-based filtering.
Of the known compact-object binaries with MS primaries (including Gaia BH 1 and several MS+NS systems), only two were included in our parent sample of 19,664 stars. The others were excluded because they lack epoch RV measurements (rv_method_used $\neq 1$) and therefore did not meet the initial selection criteria. For the two MS+NS systems that were included (Gaia DR3 5136025521527939072 and 1801110822095134848, with periods of 537 and 894 days, respectively; El-Badry et al. 2024b), our fits recovered long periods ($P_{\mathrm{fit}}\approx 600$ and 800 d) and high companion masses ($M_{\mathrm{CO,fit}}\approx 2.4$ and 2.9 M). As expected for NS companions, these fall below the $3\,M_{\odot}$ BH threshold, but would be retained if the threshold were lowered slightly.
We also cross-matched our MS sample with the catalogue of compact-object binaries from Andrew et al. (2022), which contains 4,641 candidates. Applying the same cuts to this sample (e.g. on parallax, absolute magnitude, high ruwe, rv_amplitude) yielded 1,341 stars in common with our initial MS sample. However, most of the Andrew et al. candidates were subsequently removed by the XP-based filtering of luminous MS binaries. Within our parent sample of 19,664 stars, only 178 overlapped with Andrew et al. (15, 56, and 107 in their gold, silver, and bronze categories). The inferred companion masses and orbital periods show very good agreement. After applying the full quality cuts described in Sect. 4.4, only 30 stars remained in common with our final cleaned MS+BH candidate catalogue.
To further validate the method, we compared our inferred periods and masses to Gaia DR3 NSS solutions for the 10,420 MS stars in the parent sample with AstroSpectroSB1, SB1, or Orbital entries. Figure 15 shows that our fits reproduce the NSS parameters with good correspondence, despite being based only on two observables (ruwe and $\sigma_{\mathrm{RV}}$). However, for very short periods ($\lesssim 50$ d), we often overpredict the orbital period. This bias likely arises from contamination by hierarchical systems, where short-period inner binaries inflate the RV scatter while long-period tertiaries increase ruwe values, leading our inference framework to favour intermediate periods.
Figure 15: Comparison between binary parameters inferred from Gaia summary diagnostics and reference values from the DR3 NSS catalogue for the MS sample. Shown are histograms for the 6195 stars (out of 19,664 fitted MS stars) with AstroSpectroSB1, SB1, or Orbital NSS solutions. The panels compare orbital periods (top) and companion masses (bottom). The two known Gaia NS binaries in this sample are highlighted in black. Although they rely on only two observables (ruwe and $\sigma_{\mathrm{RV}}$), our fits reproduce the catalogued parameters with good overall correspondence.